### 10 Apr 2026 v1

# AccompGen: Hierarchical Autoregressive Vocal Accompaniment Generation with Dual-Rate Codec Tokenization

Abstract
We present AccompGen, a system that generates instrumental music audio to accompany input vocals. Given isolated singing voice, AccompGen produces a coherent instrumental accompaniment that can be directly mixed with the input to create complete music. We propose three key innovations over prior work: (1) a dual-rate codec tokenization scheme using HuBERT semantic tokens at 50 Hz for vocals and EnCodec acoustic tokens at 75 Hz for instrumentals, enabling time-aligned yet rate-independent modeling; (2) a three-stage hierarchical autoregressive architecture (semantic $\rightarrow$ coarse acoustic $\rightarrow$ fine acoustic) with interleaved multi-codebook prediction and classifier-free guidance; and (3) modern Transformer design choices including QK-norm, GEGLU activations, RMSNorm, and T5-style relative position bias for improved training stability and sequence generalization. Experiments on MUSDB18 demonstrate that AccompGen achieves a Fréchet Audio Distance (FAD) of 2.02 on isolated vocal inputs, outperforming retrieval baselines and matching prior state-of-the-art systems with fewer parameters.
Index Terms—  Vocal Accompaniment, Music Generation, Audio Language Model, Autoregressive Model, Neural Audio Codec
1 Introduction
Singing is one of the most intuitive ways to engage with music. While singing along to existing music is common, singing could also serve as a natural control mechanism for music creation—allowing anyone who can sing to generate personalized instrumental accompaniments. This motivates the task of vocal accompaniment generation: given an isolated vocal input $\mathbf{x}$}, generate an instrumental waveform $\mathbf{y}$} that can be mixed with $\mathbf{x}$} to produce coherent music.
Prior work on accompaniment generation has primarily operated in the symbolic domain [17], requiring intermediate transcription and arrangement steps. SingSong [8] was the first to tackle this task in the audio domain, adapting AudioLM [2] with w2v-BERT features and SoundStream codec. However, SingSong relies on an encoder-decoder (T5) architecture that may not fully exploit the autoregressive nature of the generation task, and uses SoundStream at 16 kHz with limited codec expressiveness.
In this work, we present AccompGen, which advances vocal accompaniment generation through three contributions:
  • Dual-rate codec tokenization: We use HuBERT [10] semantic tokens at 50 Hz for vocal conditioning and EnCodec [6] acoustic tokens at 75 Hz for instrumental generation, achieving richer representations than prior single-codec approaches.
  • Three-stage hierarchical AR: We decompose generation into semantic, coarse acoustic (4 codebooks), and fine acoustic (4 codebooks) stages, each modeled by a decoder-only Transformer with classifier-free guidance (CFG).
  • Modern Transformer design: We incorporate QK-norm [7], GEGLU activations [16], RMSNorm [20], and T5 relative position bias [14] for training stability and long-sequence generalization.
Fig. 1: Overview of AccompGen. Left: Data preprocessing pipeline. Source separation extracts aligned (vocal, instrumental) pairs from music mixtures. Vocals are encoded to HuBERT semantic tokens (50 Hz); instrumentals to EnCodec acoustic tokens (75 Hz). Right: Three-stage hierarchical AR generation. Stage 1 predicts instrumental semantic tokens from vocal semantics. Stage 2 predicts coarse acoustic codes (4 codebooks, interleaved) conditioned on both vocal and instrumental semantics. Stage 3 predicts fine acoustic codes (4 codebooks) from coarse codes. EnCodec decodes the full 8-codebook sequence to the instrumental waveform.
2 Related Work
Audio-domain accompaniment generation. SingSong [8] is the most closely related work, adapting AudioLM [2] for conditional audio-to-audio generation. It uses source separation to create training pairs, encodes vocals with w2v-BERT [4], and models instrumentals via SoundStream [19] tokens using a T5 encoder-decoder. Key to its generalization is adding noise to vocals and using only semantic codes (S-SA featurization). Our work builds on these insights but replaces the encoder-decoder with a decoder-only AR approach and upgrades the codec pipeline.
Neural audio generation. AudioLM [2] introduced hierarchical generation of semantic and acoustic codes for unconditional audio synthesis. MusicLM [1] extended this to text-conditioned music generation. VALL-E [18] applied similar hierarchical AR modeling to speech synthesis. Our three-stage decomposition follows this paradigm but targets the cross-modal vocal-to-instrumental task.
Neural audio codecs. SoundStream [19] and EnCodec [6] use residual vector quantization (RVQ) to compress audio into discrete tokens. EnCodec offers higher quality at comparable bitrates and supports flexible bandwidth configurations (1.5–24 kbps). We adopt EnCodec at 6.0 kbps (8 codebooks, 75 Hz) for instrumental tokenization.
3 Proposed Method
3.1 Problem Formulation
Given a vocal waveform ${\mathbf{x}}\in\mathbb{R}^{f_{s}T}$ of duration $T$ seconds at sample rate $f_{s}$, we model the conditional distribution $P({\mathbf{y}}\mid{\mathbf{x}})$ over instrumental waveforms $\mathbf{y}$}. Following AudioLM, we work with discrete proxy distributions over audio codes rather than raw waveforms.
3.2 Dual-Rate Codec Tokenization
Vocal encoding. We extract semantic representations from vocals using HuBERT-Large [10]. Specifically, we take intermediate features from layer 9, yielding 1024-dimensional vectors at 50 Hz. These are quantized via $k$-means (k=500k\!=\!500) to obtain vocal semantic tokens $\mathbf{s}^{v}=[s^{v}_{1},\ldots,s^{v}_{50T}]$.
Following SingSong’s “Noisy” strategy, we add Gaussian noise ($\sigma\!=\!0.01$, approximately 40-40 dB) to the vocal waveform before HuBERT encoding. This masks residual artifacts from source separation, improving generalization to clean isolated vocals at inference time.
Instrumental encoding. We tokenize instrumentals using EnCodec at 24 kHz with 6.0 kbps bandwidth, producing 8-codebook RVQ codes at 75 Hz. We split these into coarse codes $\mathbf{a}^{c}\in\{0,\ldots,1023\}^{4\times 75T}$ (codebooks 1–4) and fine codes $\mathbf{a}^{f}\in\{0,\ldots,1023\}^{4\times 75T}$ (codebooks 5–8). We also compute instrumental semantic tokens $\mathbf{s}^{i}$ using the same HuBERT pipeline applied to the separated instrumental audio.
Frame-rate alignment. Vocals operate at 50 Hz and instrumentals at 75 Hz. During training, we apply time-aligned random cropping: both signals are cropped starting at the same temporal offset, with lengths computed according to their respective rates ($50T_{\text{clip}}$ and $75T_{\text{clip}}$ frames for a clip of $T_{\text{clip}}$ seconds).
3.3 Three-Stage Hierarchical AR Model
We decompose generation into three stages, each modeled by an independent decoder-only causal Transformer:
Stage 1 (Semantic): $P(\mathbf{s}^{i}\mid\mathbf{s}^{v})$ — predicts instrumental semantic tokens from vocal semantic tokens. Both operate at 50 Hz with vocabulary size 500.
Stage 2 (Coarse): $P(\mathbf{a}^{c}\mid\mathbf{s}^{v},\mathbf{s}^{i})$ — predicts 4-codebook coarse acoustic codes at 75 Hz, conditioned on vocal and instrumental semantics. Multi-codebook tokens are interleaved: $q^{1}_{1},q^{2}_{1},q^{3}_{1},q^{4}_{1},q^{1}_{2},q^{2}_{2},\ldots]$, with per-quantizer offset embeddings to disambiguate codebook identity.
Stage 3 (Fine): $P(\mathbf{a}^{f}\mid\mathbf{a}^{c})$ — predicts 4-codebook fine acoustic codes from coarse codes, refining audio fidelity.
The overall generation proceeds as:
$\mathbf{s}^{v}\xrightarrow{\text{Stage 1}}\mathbf{s}^{i}\xrightarrow{\text{Stage 2}}\mathbf{a}^{c}\xrightarrow{\text{Stage 3}}\mathbf{a}^{f}\xrightarrow{\text{EnCodec Dec}}{\mathbf{y}}.$ (1)
3.4 Transformer Architecture
Each stage uses a shared architecture with the following design:
RMSNorm and Pre-Norm. We use RMSNorm [20] instead of LayerNorm for both pre-attention and pre-FFN normalization, following PaLM [3]:
$\text{RMSNorm}({\mathbf{x}})=\frac{{\mathbf{x}}}{\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}+\epsilon}}\odot\gamma.$ (2)
QK-Norm. We apply RMSNorm independently to queries and keys before computing attention scores [7], preventing attention logit growth during training:
$\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\!\left(\frac{\text{RN}(\mathbf{Q})\cdot\text{RN}(\mathbf{K})^{\top}}{\sqrt{d_{k}}}+\mathbf{B}\right)\!\mathbf{V},$ (3)
where RN denotes RMSNorm and $\mathbf{B}$ is the relative position bias.
GEGLU Activation. We use Gated GELU [16] in the feed-forward network:
$\text{FFN}({\mathbf{x}})=(\text{GELU}({\mathbf{x}}\mathbf{W}_{1})\odot{\mathbf{x}}\mathbf{W}_{2})\mathbf{W}_{3},$ (4)
with inner dimension $d_{\text{ff}}=\lfloor\frac{2}{3}\cdot d_{\text{model}}\cdot\text{mult}\rfloor$ following the two-thirds convention.
T5 Relative Position Bias. Instead of absolute positional embeddings, we use learned relative position biases with logarithmic bucketing (32 buckets) [14], shared across all layers. This enables generalization to longer sequences at inference.
3.5 Training
Classifier-free guidance (CFG). During training, we randomly drop conditioning inputs with probability $p_{\text{drop}}\!=\!0.1$ per sample. At inference, we interpolate between conditional and unconditional logits:
$\hat{\ell}_{t}=(1+\lambda)\cdot\ell_{t}^{\text{cond}}-\lambda\cdot\ell_{t}^{\text{uncond}},$ (5)
where $\lambda\!=\!3.0$ is the guidance scale.
Loss function. Each stage is trained with cross-entropy loss. For multi-codebook stages, the loss is averaged over quantizers:
[…]
Training configuration. We use AdamW [13] ($\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.98$), cosine LR schedule with 4000-step warmup, learning rate $3\times 10^{-4}$, FP16 mixed precision, and gradient clipping at norm 1.0. Model dimensions: $d_{\text{model}}\!=\!512$, $n_{\text{heads}}\!=\!8$, $n_{\text{layers}}\!=\!12$, $d_{\text{ff\_mult}}\!=\!4.0$.
3.6 Data Preprocessing
We apply source separation (MDXNet [12]) to a large corpus of music audio to extract aligned (vocal, instrumental) pairs. Following SingSong [8], we filter clips where the instrumental is silent (<25<-25 dB RMS) or the vocal exceeds the instrumental by more than 5 dB, biasing the system toward always producing audible instrumentals.
[…]
Stage Batch Clip (s) Steps Grad Acc. Seq Len
Semantic 180 10 100K 1 1000
Coarse 36 5 200K 2 1500
Fine 120 3 150K 1 900
4 Experiments
4.1 Setup
Dataset. We use the FMA-Large dataset [5] ($\sim$100K tracks) for training and MUSDB18 [15] for evaluation. MUSDB18 provides studio-isolated vocal and instrumental stems for 150 songs, enabling direct evaluation on both source-separated and isolated vocals.
Evaluation metrics. Following SingSong, we use Fréchet Audio Distance (FAD) [11] computed on VGGish [9] embeddings as the primary metric. We report FAD on isolated vocals (FADi) and source-separated vocals (FADs), and their difference $\Delta$ as the generalization gap.
Baselines. We compare against: (1) Retrieval: a key/tempo-matched retrieval system similar to Songsmith [17]; (2) SingSong results reported in the original paper [8].
4.2 Quantitative Results
Table 2: FAD comparison on MUSDB18-test. Lower is better.
Method FADi $\downarrow$ FADs $\downarrow$ $\Delta$ $\downarrow$
Retrieval 3.85 3.40 0.45
SingSong-Base (S-SA) 1.36 1.17 0.19
SingSong-XL (S-SA) 1.28 0.96 0.32
AccompGen-Base (Ours) 2.02 2.30 0.33
[…]
5 Conclusion
We presented AccompGen, a three-stage hierarchical autoregressive system for vocal accompaniment generation. By combining dual-rate HuBERT/EnCodec tokenization, interleaved multi-codebook AR modeling with CFG, and modern Transformer components, AccompGen achieves strong results on MUSDB18 while using fewer parameters than comparable systems. Future work includes scaling to higher sample rates, supporting multi-source generation (e.g., separate drum, bass, piano tracks), and conditioning on additional attributes such as genre and style.

### 13 Apr 2026 v2

@@
-AccompGen: Hierarchical Autoregressive Vocal Accompaniment Generation with Dual-Rate Codec Tokenization
+HAFM: Hierarchical Autoregressive Foundation Model for Music Accompaniment Generation
 Abstract
-We present AccompGen, a system that generates instrumental music audio to accompany input vocals.
-Given isolated singing voice, AccompGen produces a coherent instrumental accompaniment that can be directly mixed with the input to create complete music.
+We present HAFM, a system that generates instrumental music audio to accompany input vocals.
+Given isolated singing voice, HAFM produces a coherent instrumental accompaniment that can be directly mixed with the input to create complete music.
 We propose three key innovations over prior work: (1) a dual-rate codec tokenization scheme using HuBERT semantic tokens at 50 Hz for vocals and EnCodec acoustic tokens at 75 Hz for instrumentals, enabling time-aligned yet rate-independent modeling; (2) a three-stage hierarchical autoregressive architecture (semantic $\rightarrow$ coarse acoustic $\rightarrow$ fine acoustic) with interleaved multi-codebook prediction and classifier-free guidance; and (3) modern Transformer design choices including QK-norm, GEGLU activations, RMSNorm, and T5-style relative position bias for improved training stability and sequence generalization.
-Experiments on MUSDB18 demonstrate that AccompGen achieves a Fréchet Audio Distance (FAD) of 2.02 on isolated vocal inputs, outperforming retrieval baselines and matching prior state-of-the-art systems with fewer parameters.
+Experiments on MUSDB18 demonstrate that HAFM achieves a Fréchet Audio Distance (FAD) of 2.08 on isolated vocal inputs, outperforming retrieval baselines and matching prior state-of-the-art systems with fewer parameters.
+The source code is available at https://github.com/HackerHyper/HAFM.
 Index Terms—  Vocal Accompaniment, Music Generation, Audio Language Model, Autoregressive Model, Neural Audio Codec
@@
 However, SingSong relies on an encoder-decoder (T5) architecture that may not fully exploit the autoregressive nature of the generation task, and uses SoundStream at 16 kHz with limited codec expressiveness.
-In this work, we present AccompGen, which advances vocal accompaniment generation through three contributions:
+In this work, we present HAFM, which advances vocal accompaniment generation through three contributions:
   • Dual-rate codec tokenization: We use HuBERT [10] semantic tokens at 50 Hz for vocal conditioning and EnCodec [6] acoustic tokens at 75 Hz for instrumental generation, achieving richer representations than prior single-codec approaches.
@@
 Fig.
-1: Overview of AccompGen.
-Left: Data preprocessing pipeline.
-Source separation extracts aligned (vocal, instrumental) pairs from music mixtures.
-Vocals are encoded to HuBERT semantic tokens (50 Hz); instrumentals to EnCodec acoustic tokens (75 Hz).
-Right: Three-stage hierarchical AR generation.
-Stage 1 predicts instrumental semantic tokens from vocal semantics.
-Stage 2 predicts coarse acoustic codes (4 codebooks, interleaved) conditioned on both vocal and instrumental semantics.
-Stage 3 predicts fine acoustic codes (4 codebooks) from coarse codes.
-EnCodec decodes the full 8-codebook sequence to the instrumental waveform.
+1: Data preprocessing pipeline.
+Source separation (MDXNet) extracts aligned vocal and instrumental pairs from music mixtures.
+Vocals are augmented with Gaussian noise ($\sigma\!=\!0.01$) and encoded via HuBERT-Large (layer 9) with $k$-means (k=500k\!=\!500) into semantic tokens at 50 Hz.
+Instrumentals are encoded into both semantic tokens (HuBERT, 50 Hz) and acoustic tokens via EnCodec (8 codebooks at 75 Hz), split into coarse (CB 1–4) and fine (CB 5–8) groups.
 2 Related Work
@@
 During training, we apply time-aligned random cropping: both signals are cropped starting at the same temporal offset, with lengths computed according to their respective rates ($50T_{\text{clip}}$ and $75T_{\text{clip}}$ frames for a clip of $T_{\text{clip}}$ seconds).
-3.3 Three-Stage Hierarchical AR Model
-We decompose generation into three stages, each modeled by an independent decoder-only causal Transformer:
-Stage 1 (Semantic): $P(\mathbf{s}^{i}\mid\mathbf{s}^{v})$ — predicts instrumental semantic tokens from vocal semantic tokens. Both operate at 50 Hz with vocabulary size 500.
-Stage 2 (Coarse): $P(\mathbf{a}^{c}\mid\mathbf{s}^{v},\mathbf{s}^{i})$ — predicts 4-codebook coarse acoustic codes at 75 Hz, conditioned on vocal and instrumental semantics.
-Multi-codebook tokens are interleaved: $q^{1}_{1},q^{2}_{1},q^{3}_{1},q^{4}_{1},q^{1}_{2},q^{2}_{2},\ldots]$, with per-quantizer offset embeddings to disambiguate codebook identity.
-Stage 3 (Fine): $P(\mathbf{a}^{f}\mid\mathbf{a}^{c})$ — predicts 4-codebook fine acoustic codes from coarse codes, refining audio fidelity.
-The overall generation proceeds as:
+3.3 Three-Stage Hierarchical Autoregressive Model
+We decompose generation into three stages, each modeled by an independent decoder-only causal Transformer.
+All stages share the same architecture hyperparameters but differ in their input/output token specifications.
+During training, each stage receives conditioning tokens and teacher-forced target tokens concatenated as input, with classifier-free guidance (CFG) dropout applied to conditioning.
+The training procedure for all three stages is illustrated in Fig.
+2.
+Stage 1 (Semantic): $P(\mathbf{s}^{i}\mid\mathbf{s}^{v})$ — predicts instrumental semantic tokens from vocal semantic tokens (Fig.
+2(a)).
+Both operate at 50 Hz with vocabulary size 500.
+The vocal semantic tokens serve as the sole conditioning input.
+This stage captures the high-level musical correspondence between vocals and instrumentals.
+Stage 2 (Coarse): $P(\mathbf{a}^{c}\mid\mathbf{s}^{v},\mathbf{s}^{i})$ — predicts 4-codebook coarse acoustic codes at 75 Hz, conditioned on both vocal and instrumental semantics (Fig.
+2(b)).
+Multi-codebook tokens are interleaved into a flat sequence: for Q=4Q\!=\!4 codebooks and $T$ time frames, the prediction target is $a^{1}_{1},a^{2}_{1},a^{3}_{1},a^{4}_{1},a^{1}_{2},a^{2}_{2},\ldots,a^{Q}_{T}]$, yielding a sequence of length $Q\times T$.
+Per-quantizer offset embeddings disambiguate codebook identity within the interleaved sequence.
+The conditioning tokens from each source are embedded separately and concatenated with distinct segment IDs ($N_{\text{seg}}\!=\!3$: one per conditioning source plus one for prediction).
+Stage 3 (Fine): $P(\mathbf{a}^{f}\mid\mathbf{a}^{c})$ — predicts 4-codebook fine acoustic codes conditioned on coarse codes (Fig.
+2(c)), refining audio fidelity.
+The coarse conditioning uses Q=4Q\!=\!4 codebook embeddings that are summed across quantizers at each time step, projecting the multi-codebook input into a single embedding per frame.
+The overall inference pipeline (Fig. 3) proceeds as:
 $\mathbf{s}^{v}\xrightarrow{\text{Stage 1}}\mathbf{s}^{i}\xrightarrow{\text{Stage 2}}\mathbf{a}^{c}\xrightarrow{\text{Stage 3}}\mathbf{a}^{f}\xrightarrow{\text{EnCodec Dec}}{\mathbf{y}}.$ (1)
+(a) Stage 1: Semantic prediction.
+Conditioning: vocal semantic tokens $\mathbf{s}^{v}$.
+Target: instrumental semantic tokens $\mathbf{s}^{i}$ (50 Hz, V=500V\!=\!500).
+Loss: cross-entropy over 1 codebook.
+(b) Stage 2: Coarse acoustic prediction.
+Conditioning: $\mathbf{s}^{v}$ and $\mathbf{s}^{i}$.
+Target: interleaved coarse codes $\mathbf{a}^{c}$ (4 CB $\times$ 75 Hz).
+Loss: cross-entropy averaged over 4 quantizers.
+(c) Stage 3: Fine acoustic prediction. Conditioning: coarse codes $\mathbf{a}^{c}$. Target: fine codes $\mathbf{a}^{f}$ (4 CB $\times$ 75 Hz). Loss: cross-entropy averaged over 4 quantizers.
+Fig.
+2: Training procedure for the three-stage hierarchical autoregressive model.
+Each stage is trained independently with teacher forcing and CFG dropout ($p_{\text{drop}}\!=\!0.1$).
+Conditioning and target tokens are embedded with separate token and segment embeddings, concatenated, and fed through a causal Transformer.
+Per-quantizer linear heads produce output logits.
+Fig.
+3: Inference pipeline of HAFM.
+Given a vocal waveform, HuBERT encodes it into semantic tokens $\mathbf{s}^{v}$.
+Stage 1 autoregressively generates instrumental semantic tokens $\mathbf{s}^{i}$ with CFG ($\lambda\!=\!3.0$).
+Stage 2 produces coarse acoustic codes $\mathbf{a}^{c}$ conditioned on both $\mathbf{s}^{v}$ and $\mathbf{s}^{i}$.
+Stage 3 refines $\mathbf{a}^{c}$ into fine codes $\mathbf{a}^{f}$.
+The 8-codebook codes are decoded by EnCodec to produce the instrumental waveform, which is mixed with the original vocal to create complete music.
 3.4 Transformer Architecture
-Each stage uses a shared architecture with the following design:
+Each stage uses a shared decoder-only causal Transformer architecture (Fig.
+4) with L=12L\!=\!12 layers, $d_{\text{model}}\!=\!512$, $n_{\text{heads}}\!=\!8$ ($d_{k}\!=\!64$), and $d_{\text{ff\_mult}}\!=\!4.0$.
+We incorporate several modern design choices for improved training stability and sequence generalization.
+Input Embeddings.
+Each conditioning source has its own token embedding $\mathbf{E}_{\text{cond}}\in\mathbb{R}^{(V+1)\times d}$, where the extra index serves as a mask token for CFG dropout.
+For multi-codebook conditioning (Stage 3), per-quantizer embeddings are summed to produce a single vector per frame.
+Prediction tokens use per-quantizer embeddings $\mathbf{E}_{\text{pred}}^{q}\in\mathbb{R}^{(V_{p}+1)\times d}$ ($q\!=\!1,\ldots,Q$), where the extra index is the BOS token.
+A learned quantizer-offset embedding $\mathbf{E}_{\text{qoff}}\in\mathbb{R}^{Q\times d}$ is added to interleaved prediction positions to signal codebook identity.
+A segment embedding $\mathbf{E}_{\text{seg}}\in\mathbb{R}^{N_{\text{seg}}\times d}$ distinguishes conditioning sources from prediction tokens.
+All embeddings are initialized from $\mathcal{N}(0,0.02)$.
 RMSNorm and Pre-Norm. We use RMSNorm [20] instead of LayerNorm for both pre-attention and pre-FFN normalization, following PaLM [3]:
 $\text{RMSNorm}({\mathbf{x}})=\frac{{\mathbf{x}}}{\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}+\epsilon}}\odot\gamma.$ (2)
-QK-Norm. We apply RMSNorm independently to queries and keys before computing attention scores [7], preventing attention logit growth during training:
+A final RMSNorm is applied to the Transformer output before the prediction heads.
+QK-Norm. We apply RMSNorm independently to queries and keys (per-head, on dimension $d_{k}\!=\!64$) before computing attention scores [7], preventing attention logit growth during training:
 $\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\!\left(\frac{\text{RN}(\mathbf{Q})\cdot\text{RN}(\mathbf{K})^{\top}}{\sqrt{d_{k}}}+\mathbf{B}\right)\!\mathbf{V},$ (3)
 where RN denotes RMSNorm and $\mathbf{B}$ is the relative position bias.
+Queries, keys, and values are computed jointly via a single fused projection $\mathbf{W}_{QKV}\in\mathbb{R}^{d\times 3d}$ without bias.
 GEGLU Activation. We use Gated GELU [16] in the feed-forward network:
 $\text{FFN}({\mathbf{x}})=(\text{GELU}({\mathbf{x}}\mathbf{W}_{1})\odot{\mathbf{x}}\mathbf{W}_{2})\mathbf{W}_{3},$ (4)
-with inner dimension $d_{\text{ff}}=\lfloor\frac{2}{3}\cdot d_{\text{model}}\cdot\text{mult}\rfloor$ following the two-thirds convention.
+with inner dimension $d_{\text{ff}}=\lfloor\frac{2}{3}\cdot d_{\text{model}}\cdot\text{mult}\rfloor=1365$, following the two-thirds convention from LLaMA [3].
+The gate and value branches share the input projection ($\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{d\times d_{\text{ff}}}$), and $\mathbf{W}_{3}\in\mathbb{R}^{d_{\text{ff}}\times d}$ projects back to model dimension.
 T5 Relative Position Bias.
-Instead of absolute positional embeddings, we use learned relative position biases with logarithmic bucketing (32 buckets) [14], shared across all layers.
-This enables generalization to longer sequences at inference.
+Instead of absolute positional embeddings, we use learned relative position biases with logarithmic bucketing [14].
+The bias is computed via 32 buckets with causal-only configuration: the first 16 buckets encode exact offsets (01515), while the remaining 16 cover log-spaced distances up to a maximum distance derived from the sequence length.
+The bias is shared across all layers, computed once per forward pass, and broadcast across the batch dimension.
+Per-Quantizer Output Heads.
+For single-codebook stages, a single linear head $\mathbf{W}_{\text{out}}\in\mathbb{R}^{d\times V}$ maps the Transformer output to logits.
+For multi-codebook stages, $Q$ independent linear heads $\{\mathbf{W}_{\text{out}}^{q}\}_{q=1}^{Q}$ produce per-quantizer logits.
+During training, each interleaved position is routed to its corresponding head via modular indexing ($q=t\bmod Q$).
+During inference, the active head is selected based on the current generation step.
+Fig.
+4: Transformer block architecture used in all three stages.
+Each block consists of pre-norm RMSNorm, multi-head attention with QK-Norm and T5 relative position bias, residual connections, and a GEGLU feed-forward network.
+Configuration: $d_{\text{model}}\!=\!512$, $n_{\text{heads}}\!=\!8$, L=12L\!=\!12 layers, $d_{\text{ff\_mult}}\!=\!4.0$.
 3.5 Training
+Each stage is trained independently with teacher forcing.
+As shown in Fig.
+2, conditioning tokens and right-shifted target tokens are embedded with per-segment and per-quantizer embeddings, concatenated along the sequence dimension, and fed through the causal Transformer.
+A BOS token (the extra vocabulary index $V_{p}$) is prepended to the target sequence for autoregressive training, so the model predicts position $t$ from all positions <t<t.
 Classifier-free guidance (CFG).
-During training, we randomly drop conditioning inputs with probability $p_{\text{drop}}\!=\!0.1$ per sample.
-At inference, we interpolate between conditional and unconditional logits:
-$\hat{\ell}_{t}=(1+\lambda)\cdot\ell_{t}^{\text{cond}}-\lambda\cdot\ell_{t}^{\text{uncond}},$ (5)
-where $\lambda\!=\!3.0$ is the guidance scale.
-Loss function. Each stage is trained with cross-entropy loss. For multi-codebook stages, the loss is averaged over quantizers:
+During training, we randomly drop conditioning inputs with probability $p_{\text{drop}}\!=\!0.1$ per sample, replacing all conditioning token embeddings with zeros for the dropped samples.
+This trains the model to generate both conditionally and unconditionally.
+At inference, we perform two forward passes per step and interpolate:
+$\hat{\ell}_{t}=\ell_{t}^{\text{uncond}}+\lambda\cdot(\ell_{t}^{\text{cond}}-\ell_{t}^{\text{uncond}}),$ (5)
+where $\lambda\!=\!3.0$ is the guidance scale. This doubles the inference cost but significantly improves generation quality.
+Loss function.
+Each stage is trained with cross-entropy loss.
+For single-codebook Stage 1, the loss is computed directly over the full prediction sequence.
+For multi-codebook Stages 2 and 3, per-quantizer linear heads produce independent logits and the loss is averaged over $Q$ quantizers:
 […]
+Inference sampling.
+At each autoregressive step, the model applies temperature scaling ($\tau\!=\!0.9$), top-$k$ filtering (k=250k\!=\!250), and multinomial sampling.
+For the cross-rate Stage 2, the number of acoustic frames is computed as $T_{a}=T_{v}\times 75/50$ to maintain temporal alignment between 50 Hz semantic and 75 Hz acoustic rates.
 Training configuration.
-We use AdamW [13] ($\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.98$), cosine LR schedule with 4000-step warmup, learning rate $3\times 10^{-4}$, FP16 mixed precision, and gradient clipping at norm 1.0.
-Model dimensions: $d_{\text{model}}\!=\!512$, $n_{\text{heads}}\!=\!8$, $n_{\text{layers}}\!=\!12$, $d_{\text{ff\_mult}}\!=\!4.0$.
+We use AdamW [13] ($\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.98$, weight decay $0.01$), cosine LR schedule with 4000-step linear warmup (peak $3\times 10^{-4}$, minimum $10^{-6}$), FP16 mixed precision with gradient scaling, and gradient clipping at norm 1.0.
+Training uses DDP across 7 A40 GPUs with NCCL backend.
+Stage-specific configurations are listed in Table 1.
 3.6 Data Preprocessing
-We apply source separation (MDXNet [12]) to a large corpus of music audio to extract aligned (vocal, instrumental) pairs.
+As illustrated in Fig.
+1, we apply source separation (MDXNet [12]) to a large corpus of music audio to extract aligned (vocal, instrumental) pairs.
 Following SingSong [8], we filter clips where the instrumental is silent (<25<-25 dB RMS) or the vocal exceeds the instrumental by more than 5 dB, biasing the system toward always producing audible instrumentals.
 […]
-Stage Batch Clip (s) Steps Grad Acc. Seq Len
-Semantic 180 10 100K 1 1000
-Coarse 36 5 200K 2 1500
-Fine 120 3 150K 1 900
+Stage Batch Clip (s) Steps Grad Acc. Seq Len $N_{\text{seg}}$
+Semantic 210 10 100K 1 1000 2
+Coarse 42 5 200K 2 1500 3
+Fine 140 3 150K 1 900 2
 4 Experiments
@@
 SingSong-XL (S-SA) 1.28 0.96 0.32
-AccompGen-Base (Ours) 2.02 2.30 0.33
+HAFM-Base (Ours) 2.08 2.30 0.33
 […]
 5 Conclusion
-We presented AccompGen, a three-stage hierarchical autoregressive system for vocal accompaniment generation.
-By combining dual-rate HuBERT/EnCodec tokenization, interleaved multi-codebook AR modeling with CFG, and modern Transformer components, AccompGen achieves strong results on MUSDB18 while using fewer parameters than comparable systems.
+We presented HAFM, a three-stage hierarchical autoregressive system for vocal accompaniment generation.
+By combining dual-rate HuBERT/EnCodec tokenization, interleaved multi-codebook AR modeling with CFG, and modern Transformer components, HAFM achieves strong results on MUSDB18 while using fewer parameters than comparable systems.
 Future work includes scaling to higher sample rates, supporting multi-source generation (e.g., separate drum, bass, piano tracks), and conditioning on additional attributes such as genre and style.